### v1

# Comparing zero-shot self-explanations with human rationales in multilingual text classification

Abstract
Instruction-tuned LLMs are able to provide an explanation about their output to users by generating self-explanations that do not require gradient computations or the application of possibly complex XAI methods. In this paper, we analyse whether this ability results in a good explanation by evaluating self-explanations in the form of input rationales with respect to their plausibility to humans as well as their faithfulness to models. For this, we apply two text classification tasks: sentiment classification and forced labour detection. Next to English, we further include Danish and Italian translations of the sentiment classification task and compare self-explanations to human annotations for all samples. To allow for direct comparisons, we also compute post-hoc feature attribution, i.e., layer-wise relevance propagation (LRP) and apply this pipeline to 4 LLMs (Llama2, Llama3, Mistral and Mixtral). Our results show that self-explanations align more closely with human annotations compared to LRP, while maintaining a comparable level of faithfulness.
Comparing zero-shot self-explanations with human rationales
in multilingual text classification
Stephanie Brandlthanks:    Equal contribution. Department of Computer Science University of Copenhagen brandl@di.ku.dk                        Oliver Eberle Machine Learning Group Technische Universität Berlin oliver.eberle@tu-berlin.de
1 Introduction
Providing model explanations in order to increase trust and transparency into their decision making has been a key motivation for the field of Explainable AI (XAI) even before LLMs have found their way into everyday life. Nowadays, LLMs are being used for tasks like creative writing, help with homework but also for advice giving and translation while providing self-generated explanations at the same time.111www.washingtonpost.com/technology/2024/08/04/chatgpt-use-real-ai-chatbot-conversations This makes it all the more crucial to understand the quality of those self-explanations, how reliable they are and whether their faithfulness to the model and their plausibility to humans compare with other, widely investigated, post-hoc XAI methods.
Figure 1: An example from the SST sentiment classification dataset. With rationale annotations by humans, generated by Llama3 and computed post-hoc with LRP.
In this paper, we evaluate self-explanations from two text classification tasks for which human rationale annotations are available: sentiment classification and forced labour detection. We instruct 4 different LLMs (Mistral, Mixtral, Llama2 and Llama3) to solve the respective task and generate explanations based on the input text in a zero-shot experiment. We compare these with rationales provided on the same samples from various annotation studies and with state-of-the-art post-hoc explanations for Transformers, calculated based on the layer-wise relevance propagation (LRP) framework Ali et al. (2022) for each model respectively. For sentiment classification, we consider two different subsets from two different annotation studies, one also including Italian and Danish translations next to the English original. We evaluate on plausibility to the human annotations and faithfulness to the model decision, two established evaluation methods in the XAI literature DeYoung et al. (2020); Jacovi and Goldberg (2020). We further carry out qualitative analyses, for instance on the distribution of POS tags among the most frequently selected tokens, the differences between languages and the ability to follow precise instructions. We see this controlled study as a first step into a deeper understanding of the reliability and quality of self-explanations and we believe that by covering four language models, three languages and two very different domains of text classification tasks, with different levels of difficulty and text length we can provide valuable insights. We release our code on GitHub to enable reproducibility and potential future research.
2 Related Work
Generated self-explanations come with both new opportunities and challenges. Prior work in self-explanations for text has focused on new evaluation strategies and model improvements. Ye and Durrett (2022) evaluate whether including self-explanations can improve model performance on in-context learning while Madsen et al. (2024) proposed several instruction-based self-consistency checks to measure faithfulness in generated explanations.
Another line of work by Wiegreffe et al. (2022) seeks to improve free text self-explanations with the help of human-written explanations that are included in the instruction. Similarly to Kunz and Kuhlmann (2024), self-explanations are evaluated on a variety of properties by the means of human annotation. They are found to be generally true, grammatical and factual Wiegreffe et al. (2022) and further selective, to contain illustrative examples and rarely subjective according to Kunz and Kuhlmann. For those two papers GPT-3 on CommonsenseQA/NLI and GPT-4 on the Alpaca dataset were analysed, respectively. In our study, we consider human rationales as the ground truth for explaining a decision, against which we compare model self-explanations and post-hoc attributions.
Recent work by Huang et al. (2023) investigates self-explanations by ChatGPT on sentiment classification for SST, comparing faithfulness of self-explanations against different features attribution methods. They experiment with different settings by swapping the order of classification and explanations within a single instruction prompt, asking the model for top-k rationale tokens or continuous token scores, but find no method that stands out in faithfulness while observing significant disagreement across explainability approaches.
Our work focuses on a direct comparison of plausibility and faithfulness, using binary rationales and comparing them to post-hoc LRP attributions, which have been shown to faithfully reflect LLM predictions Ali et al. (2022); Achtibat et al. (2024). Additionally, we extend our analysis to a more complex text classification task–forced labor detection–and to multilingual settings, further broadening the scope and applicability of our study.
3 Experimental Setup
3.1 Datasets
We selected two text classification datasets for sentiment analysis and forced labor detection, for which human rationale annotations have been collected. With those two dataset we cover different aspects and levels of difficulty in both classification and rationale annotation. SST has been widely used for binary sentiment classification, with rationales available in English, Italian and Danish subsets. Texts are rather short and language models have been shown to solve this task successfully, while the second dataset of longer news articles on forced labour detection is more challenging for both classification and rationale extraction, and is also less likely to have been part of the models’ pre-training.
SST We use two different subsets from the Stanford Sentiment Treebank (SST2, Socher et al. 2013) for binary sentiment classification on movie reviews. The first subset contains 263 samples from the validation and test split from SST2 with an average sentence length of 18 tokens. Human rationale annotations have been published for that subset by Thorn Jakobsen et al. (2023) where each sample has been annotated by multiple annotators, 8 on average, who were recruited via Prolific. Annotators were first asked to classify the sample into one of three classes: positive, neutral or negative where none of the sentences was assigned neutral as a gold label. In a second step, annotators should choose the parts of the input that support their label choice. We select the rationale annotations with the correct labels from the first step for further analysis. We averaged the binary rationales across all annotators (with correct label classification) and set a threshold of $0.5$ (after averaging) for the token selection. We additionally analyse the rationale annotations collected by Jørgensen et al. (2022) on a subset of 250 samples from the validation set of SST2. All samples were translated into Danish and Italian with an average sentence length of 15-17. Rationale annotation was carried out by 2 annotators per language (including English), who were native speakers with linguistic training. In contrast to the annotations collected by Thorn Jakobsen et al., the correct sentiment (positive or negative) was provided and the annotators were asked to select parts of the input that supported the gold label.
RaFoLa The authors of Mendez Guzman et al. (2022) published a Rationale-annotated corpus for Forced-Labour detection. This multi-class and multi-label dataset contains 989 English news articles that have been labelled and annotated according to 11 risk indicators defined by the International Labour Organization. Rationale annotations were carried out by two annotators who selected parts of the input to justify their label decision if they found evidence for any of the 11 labels. A subset of 100 articles was annotated by both annotators with a label agreement of $0.81$ (micro F1) and a rationale agreement of $0.73$ (intersection-over-union). The remaining articles were only annotated by one of the annotators. Each news article was assigned $1.2$ labels on average while $43\%$ were assigned with at least one label. For our analysis, we selected the 4 most frequent classes with occurrences between 117-256 out of the 989 articles. As we carry out zero-shot experiments on models that have not been fine-tuned on this task, we further convert this task into a binary classification task where we ask for a specific label once at a time. We are providing the definition of the respective forced labour indicator as part of the instruction, see Figure 9 & 10.
3.2 Rationale Extraction
For our experiments, we apply the same pipeline to the following 4 instruction fine-tuned LLMs: Llama2-13B,222meta-llama/Llama-2-13b-chat-hf Llama3.1-8B333meta-llama/Meta-Llama-3.1-8B-Instruct, Mistral-7B444mistralai/Mistral-7B-Instruct-v0.3 and quantized Mixtral-8x7B555mistralai/Mixtral-8x7B-Instruct-v0.1.
In a first step, we ask the model to classify the given text into positive/negative for SST and into yes/no depending on evidence for a specific risk indicator for the RaFoLa dataset. If the model manages to generate the correct answer, we ask it to generate rationales based on the relevant provided context of the input. In case of RaFoLa, we follow the original data collection and only request rationales if the respective risk indicator is present.
For the subsets in Italian and Danish, we have manually translated the prompts to the respective language with the help of native speakers.
3.3 LRP Relevances
To extract input attribution scores, we use layer-wise relevance propagation (LRP); a widely used and state-of-the-art XAI method to compute feature attributions Ali et al. (2022); Achtibat et al. (2024). Following the proposed propagation rules for Transformer models, we compute relevance scores for Llama and Mistral by backpropagating the logit for the correctly generated class token. While self-explanations and human annotations provide binary rationales, LRP assigns continuous scores to each token. To allow direct comparison, we used the number of tokens annotated by humans as the cut-off for LRP relevance scores in both tasks. Tokens were ranked by their scores, and the top-k tokens–where $k$ is the number of rationales from the human annotation–were selected for further analysis.
3.4 Constraining Self-Explanations
Initial experiments showed that without precise instructions, the model would return $80\%$ of the input tokens as rationales for SST where humans had annotated approximately $30\%$. This made comparisons difficult, so we chose to request a maximum number of tokens based on the number of annotated tokens by humans for each sample. Language models did not always follow this request but we could reduce the ratio of tokens to a comparable level with human annotations. For RaFoLa, this problem did not occur on the same level since the input texts were much longer and humans annotated entire phrases. We thus decided not to include an upper bound for the RaFoLa rationales. We will discuss ratios and instruction following in more details.
4 Results
We first show and discuss the main results of pair-wise agreement between the different kinds of rationales, i.e., plausibility scores, before further analysing their faithfulness, the differences between languages and a qualitative analysis.
SST
mSST      English
Danish
Italian
Figure 2: Pair-wise comparison scores (Cohen’s Kappa) between rationales on SST and multilingual SST (English, Danish and Italian). We compare human-annotated, model-generated, and post-hoc LRP rationales.
RaFoLa       #1 Abuse of vulnerability
asdf
#2 Abusive working and living conditions
#5 Excessive Overtime
asdf
#8 Physical and sexual violence
asdf
Figure 3: Pair-wise scores (Cohen’s Kappa) between rationales on RaFoLa for 4 different forced labour indicators.
4.1 Plausibility
We show pair-wise comparisons between human-annotated, model-generated and post-hoc computed rationales by calculating sample-wise Cohen’s Kappa scores between the binary scores and averaging across samples for different models. When considering human annotations as the ground truth, this is usually referred to as plausibility DeYoung et al. (2020). We here also show the comparison between post-hoc and self-generated rationales, i.e., self-explanations.
Cohen’s Kappa Cohen (1960) is a well-established method to measure inter-annotator agreement (IAA) between two annotators, in our case those are either the averaged human annotations or the two different types of model rationales (generated and post-hoc). We chose Kappa over F1 scores, which is also often used to evaluate IAA but comes with two obstacles. It is (i) driven by the imbalance of classes (here selected and not selected tokens) leading to a higher offset for annotations with a ratio of selected tokens closer to 0.5 and it (ii) does not consider randomness as a confounding factor. Both of these obstacles are taken into account when computing Cohen’s Kappa scores, leading to overall lower but more robust scores in comparison to F1. We report F1 scores in the Appendix in Section B.
[…]
4.2 Task accuracy and rationale ratios
Table 1 presents model accuracies and ratios of identified rationale tokens for SST, multilingual SST (mSST), and RaFoLa datasets in humans and models. Accuracies for SST and mSST are generally high across models, while RaFoLa shows more variation, particularly lower accuracies and rationale ratios. The ratio of identified rationale tokens by humans tends to be higher for the shorter movie reviews in SST and mSST, and lower for the longer paragraphs in RaFoLa, suggesting that the human annotators have more sparsely identified relevant phrases when presented with longer texts. This can be attributed to the differences in samples across the datasets. While movie reviews in SST tend to be short and were written with the purpose of expressing a sentiment, the articles in RaFoLa were more descriptive of a specific situation or incident and might or might not include the violation of one (or more) of the 11 risk indicators. We further can assume that most models nowadays have seen the original English version of SST during training and are thus more familiar with this type of data.
Table 1: Model accuracies for SST, multilingual SST and RaFoLa, alongside the ratio of identified rationale tokens by humans and those of self-generated rationales across models.
acc ratio ratio model
llama llama3 mistral mixtral human llama llama3 mistral mixtral
SST 0.88 0.98 0.98 0.98 0.29 0.22 0.25 0.32 0.14
mSST (EN) 0.85 0.98 0.98 1.00 0.33 0.23 0.27 0.32 0.15
mSST (DK) 0.84 0.87 0.97 0.97 0.38 0.18 0.29 0.40 0.21
mSST (IT) 0.88 0.95 0.97 0.99 0.37 0.26 0.31 0.39 0.25
RaFoLA #1 0.50 0.47 0.65 0.39 0.05 0.05 0.23 0.12 0.16
RaFoLA #2 0.58 0.58 0.58 0.48 0.06 0.04 0.21 0.11 0.12
RaFoLA #5 0.88 0.92 0.89 0.81 0.04 0.02 0.08 0.06 0.05
RaFoLA #8 0.90 0.90 0.90 0.90 0.07 0.04 0.12 0.08 0.10
4.3 Faithfulness
Besides plausibility, faithfulness is the most commonly used evaluation approach to judge the quality of model explanations. Especially for feature attribution approaches, removal of most relevant features has been used to assess how faithful a feature subset is with respect to the model prediction, i.e. if removing a highly relevant subset will lead to a strong decrease of the prediction. We evaluated faithfulness by measuring the change in probability after masking the tokens as identified by the different rationales (human, self-generated and post-hoc).
Compared to human and self-generated rationales, post-hoc attributions provide a relevance score for each token in the input prompt, requiring to binarize post-hoc attributions to allow for direct comparison as described in Section 3.3. Additionally, we included a baseline that randomly removes as many token as identified by humans. Our results for SST and RaFoLa are summarized in Figure 4. Average initial probability for the correct answer token is given as a dashed line for each model.
We find that levels of faithfulness for humans, generated, and post-hoc explanations are comparable, with human explanations being overall as faithful as those generated by models or post-hoc methods. For some cases, e.g. class #5 for RaFoLA, human rationales provide more faithful feature attributions than mistral explanations.
Self-generated model explanations can be more faithful than post-hoc ones, and the reverse case can also be true, depending on the task and model. Overall, they provide similarly faithful model rationales. We further investigated the limited impact of removing features on the class token probability in the post-hoc setting and found that the most relevant tokens are typically part of the provided task instruction, such as the class definition or question. Contrastive explanations have been proposed to provide more task-specific attributions Yin and Neubig (2022) and we provide an additional analysis of the plausbility of post-hoc explanations in Section 4.8.
SST
mSST      English
Danish
Italian
RaFoLa        #1
#2
#5
#8
Figure 4: Faithfulness evaluation for SST and mSST (top row) and RaFoLa (bottom row). The probability after masking the tokens extracted from human rationale annotations, self-generated model rationales and post-hoc model attributions is compared across models. Lower probability indicates more faithful identification of rationales.
4.4 Languages
From the set of models we consider for this study, all have been pre-trained on English data while Llama3 and Mixtral also have been pre-trained on Italian, but none of the model officially supports Danish. Considering this difference in language exposure, our results show that all models are able to solve the sentiment classification task (see Table 1) and are also able to extract meaningful rationales with plausibility on the same (or even superior) level than English. Plausibility scores between self-explanations and humans are ranging $0.46-0.74$ for Danish and Italian versus $0.4-0.69$ for English. Llama2 shows the biggest gap between English ($0.67$) and Danish/Italian ($0.46/0.52$) while Mixtral shows an increase in plausibility from English ($0.4$) to ($0.57$) for Danish. Llama3 shows high and stable plausibility scores across all languages ($\sim 0.7$). It seems all models can handle both Danish and Italian, whether this is based on data contamination during pre-training, e.g., training data for Llama2 has been reported to contain $0.11\%$ Italian Touvron et al. (2023), is unclear. Previous work has focused on zero-shot abilities in unseen languages but has mostly followed approaches with either multilingual models or translating test sets into English Ebrahimi and von der Wense (2024); Artetxe et al. (2023). To the best of our knowledge, there has not been any published work in a truly zero-shot fashion, we will leave that for future work.
Table 2: Most frequent tokens in the RaFoLa corpus(first row) and in rationales identified by human annotators, as well as self-generated and post-hoc explanations of Llama3 and Mistral for #1 and #8.
#1 Abuse of vulnerability #8 Physical and sexual violence
corpus said, workers, labour, human, work, forced, rights, slavery
human workers, work, forced, women, children, labour, said, exploitation sexual, abuse, harassment, women, violence, said, verbal, physical
llama3 workers, work, said, labour, forced, women, children, working said, women, sexual, abuse, harassment, workers, violence, physical
llama3 post-hoc labour, said, slavery, vulnerable, workers, according, trafficking, forced violence, said, harassment, abuse, report, based, sexual, physical
mistral workers, work, forced, labour, children, women, working, conditions sexual, women, harassment, said, abuse, physical, children, workers
mistral post-hoc said, trafficking, forced, labour, slavery, abuse, workers, 2020 sexual, violence, said, abuse, harassment, based, report, women
4.5 Selected tokens
[…]
4.6 POS analysis
Figure 5: Top-6 POS tags of extracted rationales from SST monolingual for the human annotated rationales (left), self-explanations by llama3 (center) and post-hoc explanations based on llama3.
In order to gain a deeper insight into the differences between rationale selection, we are extracting part-of-speech (POS) tags of each selected tokens. We show results for the monolingual SST dataset in Figure 5. We see that for humans and self-explanations adjectives are most common, closely followed by nouns, verbs and with a bigger gap adverbs. For the post-hoc rationales, nouns were most often selected followed by verbs and adjectives. This means, the top-3 POS tags are the same across the different selection methods but differ in ranking between post-hoc explanations and the others. We see similar effects across models and languages, i.e., there is higher agreement in POS rankings between humans and self-explanations in comparison to post-hoc rationales. There are a few variations in ranking across languages, which has been seen in eye-tracking fixation times and attention scores before Brandl and Hollenstein (2022).
llama2 llama3 mistral mixtral
#tokens 0.11 0.04 0.34 0.16
json-syntax 0.86 0.02 0.05 0.37
Table 3: Ratio of mismatches when prompting models for a max. number of tokens and correct json syntax. Averaged scores for SST across all subsets & languages.
4.7 Instruction following
For generating and processing the self-explanations, we instructed the models to return rationales in a json format. Since many of those outputs resulted in SyntaxErrors, we included a syntax check based on Llama3 where we instructed the model in a separate step to correct the json syntax in case such an error occurred. The ability to return correct syntax varied across models. We also saw differences when following the instruction of returning the correct number of rationales for which we set an upper bound for SST based on the human annotations. We show results for SST for all 4 models, averaged across subsets and languages in Table 3. The results show that Llama2 has a lot of difficulties with respect to json syntax with syntax errors occurring in $86\%$ of the case. Both Llama3 and Mistral have a low error rate with $2\%$ and $5\%$ respectively. At the same time, Mistral returns more than the maximum number of requested rationale tokens in 1 out of 3 instructions where Llama3 follows the instruction in most cases. Analysing and evaluating the ability to follow instructions has previously been discussed in Qin et al. (2024); Zeng et al. (2024).
4.8 Contrastive LRP
SST Refer to caption
RaFoLa Refer to caption
Figure 6: Comparing plausibility scores for non-contrastive and contrastive post-hoc approaches using Kappa agreement scores. Left: SST and multilingual SST for English, Danish and Italian (top to bottom rows). Right: RaFoLa for classes #1, #2, #5, #8 (cf. Figure 3).
In a complementary analysis, we test the alignment of model rationales with post-hoc attributions, examining whether there is a difference in the plausibility of contrastive and non-contrastive post-hoc explanations. Prior work has suggested that contrastive explanations are more aligned with human reasoning and are thus considered more valuable for humans to understand the model’s decisions Lipton (1990); Miller (2019); Jacovi et al. (2021).
Comparing Cohen’s Kappa scores shown in Figure 6 suggests that contrastive post-hoc approaches do not generally result in higher plausibility than non-contrastive ones. Similarly, contrastive explanations overall exhibit a similar level of faithfulness as non-contrastive explanations (cf. Figure 4). Although contrastive explanations can be more faithful and plausible in certain cases, such as SST and mSST (English) for Llama3, there is no consistent difference between the two approaches. This aligns with previously reported high correlation between them Eberle et al. (2023) and the minor observed performance differences for generating the correct label based on contrastive or non-contrastive post-hoc approaches Krishna et al. (2023).
5 Discussion
In this paper, we investigated explanations generated by 4 different instruction-tuned LLMs for 2 text classification tasks in English but also in Italian and Danish. Those generated explanations, i.e., self-explanations, were limited to the input tokens of the respective text samples. We compared those self-explanations on the tasks of sentiment classification (SST) and forced labour classification (RaFoLa) with human annotations for the same samples and post-hoc feature attribution with layer-wise relevance propagation (LRP). Pairwise comparison between the three different types of explanations (humans, generated, post-hoc) shows that human annotations and generated explanations agree on a much higher level than post-hoc with any of them. We further saw that Llama3 has shown the highest level of agreement across both tasks and all languages and followed the instructions more closely in comparison to the other models. Our POS analysis confirmed this finding and showed higher agreement with the “type” of tokens selected by humans compared to post-hoc rationales.
Besides established post-hoc attribution approaches, the ability of language models to provide self-generated explanations, i.e., self-explanations, has offered a direct and human-understandable communication between user and model. This not only enhances usability in particular for lay people but, as presented here, also results in explanations that are similarly faithful and align more closely with human rationales compared to binarized post-hoc attributions. Herein, the generation process behind self-explanations remains obfuscated and may suffer from counterfactuality, enabling the model to give untruthful explanations for correct predictions Ji et al. (2023). We further find that current language models require careful instructions to provide useful self-generated explanations.
While human plausibility is a desired property of predictions and explanations made by machine learning systems, LLMs are capable of identifying alternative solutions to task-solving that may not be intuitive to humans. Good explanations should thus also highlight the learned prediction strategy that is faithful to the one used by the model–even if it is not directly plausible Agarwal et al. (2024). In contrast to post-hoc attributions, which are directly derived from the prediction score (e.g., using explanatory gradients as for the LRP attributions considered here), the self-generation process is not explicitly tied to the prediction in a clear manner.
Attribution approaches herein consider the entire input to rank relevant features, including both the provided context and all instructions. Instruction templates are crucial for understanding and solving tasks, and we find that tokens related to task instructions often play a more significant role than the context itself. While this effectively identifies relevant task features from the perspective of the model, users are typically more interested in understanding which parts of the provided evidence are most relevant. Disentangling task instructions from contextual evidence is essential for aligning insights with user expectations and should inform the development of future interpretability methods.
Our study represents a first step toward understanding and building more intuitive model explanations by directly comparing human annotations with those generated by models. Evaluating free-text explanations for factuality, usability, and faithfulness is crucial for ensuring their practical and intuitive application, especially given the growing use of increasingly complex LLMs by lay people who may not fully understand their mechanisms
We currently do not know why self-explanations align more closely with human annotations than post-hoc attributions. While training procedures such as reinforcement learning from human feedback Ziegler et al. (2019) may incentivize more human-like explanations Agarwal et al. (2024), limited access to models, procedures, and datasets, restricts detailed analysis.
The ability of the models to solve the sentiment classification task in unseen languages like Danish is remarkable. Not only is the model able to understand the Danish prompt, it also is able to return specific input tokens that are aligned with the human selection. Future work could investigate whether this zero-shot performance is specific to certain language families and tasks or if it extends to more complex instructions and datasets.
Conclusion As language models have advanced and become widely adopted, understanding the mechanisms behind their predictions has emerged as a critical challenge. Using human rationale annotations across three datasets, we have investigated the human plausibility and faithfulness of self-generated explanations to post-hoc feature attributions. We find that self-generated explanations agree more closely with human rationales than post-hoc feature attribution, while remaining faithful to the model prediction.
Limitations
We acknowledge that annotations may be affected by annotator bias, varying guidelines, and differing expertise, impacting the consistency of rationales. Also the number of annotators and the level of details in the instructions varied across the annotation studies we have considered for this paper. Furthermore, for the forced labour detection, annotations by legal scholars might differ from the ones provided and would also be interesting to compare with model rationales.
We focus our study on rationales based on the input while free text explanations might provide more useful information and pose the more realistic scenario.
While agreement between human and model rationales may be desired, it has been shown in previous work, that humans do not necessarily prefer human-written explanations in comparison to the ones generated by LLMs in the case of free text explanations Wiegreffe et al. (2022).
The high zero-shot performance, especially with SST, may be an effect of data contamination, which is likely part of the training data. We can further not exclude the possibility that rationales or task explanations have been included in the training corpus.
Acknowledgments
We thank our colleagues at the CoAStaL NLP group for constructive feedback in the beginning of the project. In particular, we would like to thank Alice Schiavone and Anders Søgaard for helping us with the prompt translations.
SB received funding by the European Union under the Grant Agreement no. 10106555, FairER. OE is funded by the German Ministry for Education and Research (under refs 01IS18056A and 01IS18025A) and BIFOLD. Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or European Research Executive Agency (REA). Neither the European Union nor REA can be held responsible for them.
Appendix A Instructions
A.1 SST
Figure 7: Prompts in all 3 languages to solve sentiment classification.
Figure 8: Follow-up prompts in all 3 languages to extract rationales.
A.2 RaFoLa
Figure 9: Prompts for classification and rationale extraction for the RaFoLa dataset.
Figure 10: Indicators defined by the International Labour Organization and published by Mendez Guzman et al..
Appendix B F1-Plausibility scores
SST
mSST      English
Danish
Italian
Figure 11: Pair-wise F1 comparison scores between rationales on SST and multilingual SST (English, Danish and Italian). We compare rationales annotated by humans, generated by models, and computed post-hoc with LRP.
RaFoLa       #1 Abuse of vulnerability
#2 Abusive working and living conditions
#5 Excessive Overtime
#8 Physical and sexual violence
Figure 12: Pair-wise F1 comparison scores between rationales on RaFoLa.
Appendix C Most frequent rationale tokens
Table 4: List of top-8 most frequent tokens in the RaFoLa corpus (first row) together with the most frequent rationales as identified by human annotators, as well as self-generated and post-hoc explanations.
#1 Abuse of vulnerability #2 Abusive working and living cond. #5 Excessive overtime #8 Physical and sexual violence
corpus said, workers, labour, human, work, forced, rights, slavery
human workers, work, forced, women, children, labour, said, exploitation workers, conditions, work, forced, water, little, said, working hours, work, week, days, long, claimed, worked, like sexual, abuse, harassment, women, violence, said, verbal, physical
llama2 workers, work, migrant, women, exploitation, forced, labour, children workers, work, conditions, forced, day, working, children, water pay, work, ahmad, received, little, breaks, \$, 600 abuse, sexual, women, harassment, retaliation, verbal, physical, advances
llama2 post-hoc trafficking, world, china, children, said, forced, child, exploitation trafficking, forced, said, world, children, conditions, labour, covid-19 covid-19, workers, thailand, thai, kingdom, california, hours, basi said, violence, women, based, walmart, global, guardian, jennifer
llama3 workers, work, said, labour, forced, women, children, working workers, work, said, conditions, working, labour, forced, day hours, working, day, work, pay, said, workers, days said, women, sexual, abuse, harassment, workers, violence, physical
llama3 post-hoc labour, said, slavery, vulnerable, workers, according, trafficking, forced said, workers, labour, according, conditions, slavery, work, trafficking hours, working, work, overtime, day, workers, forced, said violence, said, harassment, abuse, report, based, sexual, physical
mistral workers, work, forced, labour, children, women, working, conditions workers, work, said, forced, conditions, working, labour, children said, day, mr, work, hours, days, delivery, home sexual, women, harassment, said, abuse, physical, children, workers
mistral post-hoc said, trafficking, forced, labour, slavery, abuse, workers, 2020 said, covid-19, 2019, workers, trafficking, labour, forced, world employer, covid-19, said, years, police, cotton, mr, paying sexual, violence, said, abuse, harassment, based, report, women
mixtral workers, work, said, forced, children, labour, women, abuse workers, work, said, conditions, forced, labour, working, paid hours, work, day, forced, workers, working, days, week sexual, women, said, workers, harassment, abuse, violence, reported
Table 5: List of top-8 most frequent tokens in the SST and mSST corpus together with the most frequent rationales as identified by human annotators, as well as self-generated and post-hoc explanations.
SST mSST English mSST Danish mSST Italian
full corpus movie, film, like, comedy, -, characters, work, romantic film, movie, characters, bad, like, performances, funny, story film, filmen, karakterer, bare, sjov, ‘, se, præstationer film, i, divertente, personaggi, interpretazioni, trama, avvincente, commedia
human funny, movie, beautifully, bad, best, hilarious, stupid, wonderful bad, performances, funny, good, dull, film, compelling, dumb film, sjov, overbevisende, plot, vittig, bare, dårligt, præstationer divertente, avvincente, noioso, interpretazioni, ben, brutto, film, intelligente
llama2 bad, beautifully, best, fun, stupid, wonderful, funny, worst bad, funny, dull, compelling, witty, good, long, little sjov, spændende, underholdende, præstationer, bedste, overbevisende, spænding, vittig divertente, film, avvincente, noioso, interpretazioni, ben, intelligente, senso
llama2 post-hoc movie, comedy, film, little, like, far, funny, stupid movie, film, bad, funny, dull, little, dumb, compelling film, filmen, ‘, sjov, præstationer, karakterer, dårlig, dårligt film, divertente, noioso, avvincente, senso, umorismo, brutto, i
llama3 bad, best, beautifully, compelling, funny, love, little, hilarious funny, bad, performances, dull, compelling, good, little, best sjov, bedste, overbevisende, humor, dårlig, dårligt, kedelig, spænding divertente, avvincente, noioso, umorismo, senso, brutto, intelligente, spiritoso
llama3 post-hoc comedy, like, far, bad, beautifully, best, little, love comedy, like, good, little, big, high, bad, dull ¦, pã¥, vã, film, sã¥, re, spã, sjov film, divertente, cosã¬, avvincente, piã¹, noioso, interpretazioni, brutto
mistral bad, funny, comedy, best, beautifully, compelling, love, far bad, performances, funny, characters, dull, compelling, good, little sjov, filmen, præstationer, film, overbevisende, sjovt, spænding, humor film, divertente, personaggi, interpretazioni, avvincente, noioso, trama, brutto
mistral post-hoc movie, film, bad, -, like, love, feels, year film, bad, funny, dull, good, dumb, characters, comedy film, filmen, filmens, sjov, ‘, præstationer, blanding, spændende film, divertente, interpretazioni, personaggi, commedia, noioso, cinema, avvincente
mixtral best, film, laugh, like, year, love, pretty, good bad, funny, compelling, great, good, witty, performances, intelligent sjov, sjovt, film, præstationer, overbevisende, humor, spændende, tilfredsstillende divertente, film, interpretazioni, noioso, umorismo, i, commedia, trama