### v2

# Comparing zero-shot self-explanations with human rationales in text classification

Abstract
Instruction-tuned LLMs are able to provide an explanation about their output to users by generating self-explanations. These do not require gradient computations or the application of possibly complex XAI methods. In this paper, we analyse whether this ability results in a good explanation. We evaluate self-explanations in the form of input rationales with respect to their plausibility to humans as well as their faithfulness to models. We study two text classification tasks: sentiment classification and forced labour detection, i.e., identifying pre-defined risk indicators of forced labour. In addition to English, we include Danish and Italian translations of the sentiment classification task and compare self-explanations to human annotations for all samples. To allow for direct comparisons, we also compute post-hoc feature attribution, i.e., layer-wise relevance propagation (LRP) and analyse 4 LLMs. We show that self-explanations align more closely with human annotations compared to LRP, while maintaining a comparable level of faithfulness. This finding suggests that self-explanations indeed provide good explanations for text classification.
Comparing zero-shot self-explanations with human rationales
in text classification
Stephanie Brandlthanks:    Equal contribution. Center for Social Data Science University of Copenhagen stephanie.brandl@sodas.ku.dk                        Oliver Eberle Machine Learning Group Technische Universität Berlin oliver.eberle@tu-berlin.de
1 Introduction
Providing model explanations to increase trust and transparency is a key motivation for the field of Explainable AI (XAI), with LLMs offering new ways to trace model decision-making. Nowadays, LLMs are being used for a wide range of tasks, ranging from creative writing and homework assistance to offering advice and translation, while providing self-generated explanations, i.e., self-explanations, in the process.111www.washingtonpost.com/technology/2024/08/04/chatgpt-use-real-ai-chatbot-conversations This makes it all the more important to understand the quality of those self-explanations, how reliable they are and whether their faithfulness to the model and their plausibility to humans compare with other, widely investigated, post-hoc XAI methods.
Figure 1: An example from the SST sentiment classification dataset. With rationale annotations by humans, generated by Llama3 and computed post-hoc with LRP.
In this paper, we evaluate self-explanations from two text classification tasks for which human rationale annotations are available: sentiment classification and forced labour detection, i.e., identifying pre-defined risk indicators of forced labour. We instruct 4 different LLMs (Mistral, Mixtral, Llama2 and Llama3) to solve the respective task and generate explanations based on the input text in a zero-shot experiment. We compare these with rationales provided on the same samples from various annotation studies and with state-of-the-art post-hoc explanations for Transformers, calculated based on the layer-wise relevance propagation (LRP) framework Ali et al. (2022) for each model respectively, see Figure 1 for an example. For sentiment classification, we consider two different subsets from two different annotation studies, one also including Italian and Danish translations alongside the English original. Following established evaluation methods in the XAI literature DeYoung et al. (2020); Jacovi and Goldberg (2020), we assess the plausibility of model rationales by measuring their agreement with human annotations and evaluate faithfulness by determining the importance of selected tokens for the model’s decision. We extend our analysis by exploring differences across rationales and task settings, comparing distributions of POS tags, named entities, and frequently selected tokens. Our study takes an initial step toward a better understanding of the reliability and quality of self-explanations, for which we analyzed four language models, three languages, and two distinct text classification domains, varying in difficulty and text length. Our findings thus provide relevant insights for model interpretability and user trust in self-explanations, and we further support reproducibility and future research by openly releasing our code.
Contributions The main contributions of this work is (i) a controlled study to compare human annotations with model explanations generated by LLMs and computed post-hoc. (ii) We evaluate plausibility to humans, i.e., the level of agreement between model and human rationales and (iii) faithfulness, i.e., the relevance of selected rationale tokens for the task (model decision). (iv) We study two different text classification tasks: sentiment classification and forced labour detection, i.e., identifying pre-defined risk indicators of forced labour. We include (v) Danish and Italian translations for the sentiment classification task alongside their English original. We further (vi) provide a qualitative analysis into the differences across languages/risk indicators with respect to frequent tokens and POS.
2 Related Work
Generated self-explanations present both new opportunities and challenges. Prior work in self-explanations for text has focused on new evaluation strategies and model improvements. Ye and Durrett (2022) evaluate whether including self-explanations can improve model performance on in-context learning while Madsen et al. (2024) proposed instruction-based self-consistency checks to measure faithfulness in generated explanations.
Another line of work by Wiegreffe et al. (2022) seeks to improve free text self-explanations with the help of human-written explanations that are included in the instruction. Similarly to Kunz and Kuhlmann (2024), self-explanations are evaluated on a variety of properties by the means of human annotation. They are found to be generally true, grammatical and factual Wiegreffe et al. (2022) and further selective, to contain illustrative examples and rarely subjective according to Kunz and Kuhlmann. For those two papers GPT-3 on CommonsenseQA/NLI and GPT-4 on the Alpaca dataset were analysed, respectively.
In our study, we consider human rationales as the ground truth for explaining a decision, against which we compare model self-explanations and post-hoc attributions.
Recent work by Huang et al. (2023) investigates self-explanations by ChatGPT on sentiment classification for SST, comparing faithfulness of self-explanations against different features attribution methods. They experiment with different settings by swapping the order of classification and explanations within a single instruction prompt, asking the model for top-k rationale tokens or continuous token scores, but find no method that stands out in faithfulness while observing significant disagreement across explainability approaches.
Our work focuses on a direct comparison of plausibility and faithfulness, using binary rationales and comparing them to post-hoc LRP attributions, which have been shown to faithfully reflect LLM predictions Ali et al. (2022); Achtibat et al. (2024). In particular, we extend our analysis beyond commonly used sentiment classification by considering more complex tasks such as forced labor detection from news articles and multilingual settings, further broadening the scope and applicability of our study.
3 Experimental Setup
3.1 Datasets
We selected two text classification datasets for sentiment analysis and forced labor detection, for which human rationale annotations have been collected. With those two dataset we cover different aspects and levels of difficulty in both classification and rationale annotation. SST has been widely used for binary sentiment classification, with rationales available in English, Italian and Danish subsets. Texts are rather short and language models have been shown to solve this task successfully, while the second dataset of longer news articles on forced labour detection is more challenging for both classification and rationale extraction, and is also less likely to have been part of the models’ pre-training.
SST/mSST We use two different subsets from the Stanford Sentiment Treebank (SST2, Socher et al. 2013) for binary sentiment classification on movie reviews. The first subset (SST) contains 263 samples from the validation and test split from SST2 with an average sentence length of 18 tokens. Human rationale annotations have been published for that subset by Thorn Jakobsen et al. (2023) where each sample has been annotated by multiple annotators, 8 on average, who were recruited via Prolific. Annotators were first asked to classify the sample into one of three classes: positive, neutral or negative where none of the sentences was assigned neutral as a gold label. In a second step, annotators should choose the parts of the input that support their label choice. We select the rationale annotations with the correct labels from the first step for further analysis. We averaged the binary rationales across all annotators (with correct label classification) and set a threshold of $0.5$ (after averaging) for the token selection. We additionally analyse the rationale annotations collected by Jørgensen et al. (2022) on a subset of 250 samples from the validation set of SST2 (mSST). All samples were translated into Danish and Italian with an average sentence length of 15-17. Rationale annotation was carried out by 2 annotators per language (including English), who were native speakers with linguistic training. In contrast to the annotations collected by Thorn Jakobsen et al., the correct sentiment (positive or negative) was provided and the annotators were asked to select parts of the input that supported the gold label.
RaFoLa The authors of Mendez Guzman et al. (2022) published a Rationale-annotated corpus for Forced-Labour detection. This multi-class and multi-label dataset contains 989 English news articles that have been labelled and annotated according to 11 risk indicators defined by the International Labour Organization. Rationale annotations were carried out by two annotators who selected parts of the input to justify their label decision if they found evidence for any of the 11 labels. A subset of 100 articles was annotated by both annotators with a label agreement of $0.81$ (micro F1) and a rationale agreement of $0.73$ (intersection-over-union). The remaining articles were only annotated by one of the annotators. Each news article was assigned $1.2$ labels on average while $43\%$ were assigned with at least one label. For our analysis, we selected the 4 most frequent classes with occurrences between 117-256 out of the 989 articles. As we carry out zero-shot experiments on models that have not been fine-tuned on this task, we further convert this task into a binary classification task where we ask for a specific label once at a time. We provide the definition of the respective forced labour indicator as part of the instruction, see Figure 8 and Figure 9.
3.2 Rationale Extraction
For our experiments, we evaluate the following 4 instruction fine-tuned LLMs: Llama2-13B, Llama3.1-8B, Mistral-7B and quantized Mixtral-8x7B. 222meta-llama/Llama-2-13b-chat-hf333meta-llama/Meta-Llama-3.1-8B-Instruct444mistralai/Mistral-7B-Instruct-v0.3555mistralai/Mixtral-8x7B-Instruct-v0.1
In a first step, we ask the model to classify the given text into positive/negative for SST and into yes/no depending on evidence for a specific risk indicator for the RaFoLa dataset. If the model manages to generate the correct answer, we ask it to generate rationales based on the relevant provided context of the input. In case of RaFoLa, we follow the original data collection and only request rationales if the respective risk indicator is present. For the subsets in Italian and Danish, we have manually translated the prompts to the respective language with the help of native speakers.
Experimental Details The experiments are based on the transformers library. We set the repetition penalty to 1.0 and adjust the maximum length of generated text with respect to the task and expected output. We ensure reproducibility of our results by consistently using the same set of three seeds across our experiments and will release our code upon publication, including all parameters and the exact libraries used. All instructions with class definitions are presented in Appendix A.
3.3 Post-hoc Attribution
To extract input attribution scores, we use layer-wise relevance propagation (LRP); a widely used and state-of-the-art XAI method to compute feature attributions in LLMs Ali et al. (2022); Achtibat et al. (2024); Rezaei Jafari et al. (2024). Following the proposed propagation rules for Transformer models, we compute relevance scores for Llama and Mistral by backpropagating the logit for the correctly generated class token. While self-explanations and human annotations provide binary rationales, LRP assigns continuous scores to each token. To allow direct comparison, tokens are ranked based on their relevance scores, and the top-$k$ tokens–where $k$ is the number of rationales from the human annotation–are selected for further analysis.
3.4 Constraining Self-Explanations
Initial experiments show that without precise instructions, the model returns $80\%$ of the input tokens as rationales for SST where humans had annotated approximately $30\%$. This made comparisons difficult, so we chose to request a maximum number of tokens based on the number of annotated tokens by humans for each sample. Language models did not always follow this request but we could reduce the ratio of tokens to a comparable level with human annotations. For RaFoLa, this issue was less pronounced, as the input texts were much longer and humans annotated entire phrases. We thus decided not to include an upper bound for the RaFoLa rationales. We will discuss ratios and instruction following in the Appendix in Section C.
Table 1: Model accuracies for SST, multilingual SST, and RaFoLa, with highest scores shown in bold.
Data Acc. llama llama3 mistral mixtral
SST 0.88 0.98 0.98 0.98
mSST (EN) 0.85 0.98 0.98 1.00
mSST (DK) 0.84 0.87 0.97 0.97
mSST (IT) 0.88 0.95 0.97 0.99
RaFoLA #1 0.50 0.47 0.65 0.39
RaFoLA #2 0.58 0.58 0.58 0.48
RaFoLA #5 0.88 0.92 0.89 0.81
RaFoLA #8 0.90 0.90 0.90 0.90
4 Main Results
We first show and discuss the main results of task accuracy and pair-wise agreement between the different types of rationales, i.e., plausibility scores, before further analysing their faithfulness.
4.1 Task accuracy
Table 1 presents task accuracies for SST, multilingual SST (mSST), and RaFoLa. Accuracies for SST and mSST are generally high across models, while RaFoLa shows more variation and overall lower accuracies. We can assume that most models nowadays have seen the original English version of SST during training and are thus more familiar with this type of data. From the set of models we consider for this study, all have been pre-trained on English data while Llama3 and Mixtral also have been pre-trained on Italian, but none of the model officially supports Danish. Considering this difference in language exposure, our results show that all models are able to solve the sentiment classification task in Danish and Italian with accuracies comparable to English.
SST
mSST      English
Danish
Italian
RaFoLa       #1 Abuse of vulnerability
asdf
#2 Abusive working and living conditions
#5 Excessive Overtime
asdf
#8 Physical and sexual violence asdf
Figure 2: Pair-wise comparison scores (Cohen’s Kappa) between rationales on SST and multilingual SST (upper) and RaFoLa (lower). We compare human-annotated, model-generated (self-explanations), and post-hoc rationales.
4.2 Plausibility
Following XAI literature, we assess plausibility of rationales to humans by considering human annotations as the ground truth and compute agreement to model rationales DeYoung et al. (2020). Here, we further include agreement scores between model-generated and post-hoc LRP rationales. We show pair-wise comparisons by calculating sample-wise Cohen’s Kappa scores between the binary scores and averaging across samples for different models.
Metric Cohen’s Kappa Cohen (1960) is a well-established method to measure inter-annotator agreement (IAA) between two annotators, in our case those are either the averaged human annotations or the two different types of model rationales (generated and post-hoc). We chose Kappa over F1 scores, which is also often used to evaluate IAA but comes with two obstacles. It is (i) driven by the imbalance of classes (here selected and not selected tokens) leading to a higher offset for annotations with a ratio of selected tokens closer to 0.5 and it (ii) does not consider randomness as a confounding factor. Cohen’s Kappa scores account for both issues, leading to overall lower but more robust scores than F1, which are also reported in the Appendix in Section B.
[…]
SST
mSST      Danish
RaFoLa        article #1
article #8
Figure 3: Faithfulness evaluation for SST, mSST (Danish) and RaFoLa (articles #1 and #8). Model probability after masking the tokens extracted from human rationales, model self-explanation rationales and post-hoc model attributions is compared across models. Lower probability indicates more faithful identification of rationales.
4.3 Faithfulness
Besides plausibility, faithfulness is the most commonly used evaluation approach to judge the quality of model explanations. Especially for feature attribution approaches, removal of most relevant features has been used to assess how faithful a feature subset is with respect to the model prediction, i.e., if removing a highly relevant subset will lead to a strong decrease of the prediction. We evaluate faithfulness by measuring the change in probability after masking the tokens as identified by the different rationales (human, model self-explanations and model post-hoc).
Compared to human and self-explanation rationales, post-hoc attributions provide a relevance score for each token in the input prompt, requiring to binarize post-hoc attributions to allow for direct comparison as described in Section 3.3. Additionally, we included a baseline that randomly removes as many token as identified by humans. Our key results for SST and RaFoLa are summarized in Figure 3, with additional results across all languages and articles presented in Figure 15 in the Appendix. Average initial probability for the correct answer token is given as a dashed line for each model.
We find that levels of faithfulness for humans, generated, and post-hoc explanations are comparable, with human explanations being overall as faithful as those generated by models or post-hoc methods. Self-generated model explanations can be more faithful than post-hoc ones, and the reverse case can also be true, depending on the task and model. Overall, they provide similarly faithful model rationales. We further investigate the limited impact of removing features on the class token probability in the post-hoc setting and found that the most relevant tokens are typically part of the provided task instruction, such as the class definition or question. Furthermore, contrastive explanations have been proposed to provide more task-specific attributions Yin and Neubig (2022), which we analyzed in Appendix E.2. We found no critical effect of contrastive explanations on the plausibility and faithfulness of model-based rationales.
[…]
Figure 4: Distribution of POS-tags in comparison to top-6 POS in human annotations for SST. Absolute POS tags have been normalized (so they sum to 1) prior to computing the difference to human baseline. Ranking is based on prevalence in human rationales. Difference is shown for model-generated, i.e., self-explanations (left) and post-hoc LRP (right) across all data splits and languages for Mistral.
#1 Abuse of vulnerability #8 Physical & sexual violence
corpus said, workers, labour, human, work, forced, rights, slavery
human workers, work, forced, women, children, labour, said, exploitation sexual, abuse, harassment, women, violence, said, verbal, physical
llama3 workers, work, said, labour, forced, women, children, working said, women, sexual, abuse, harassment, workers, violence, physical
llama3 post-hoc labour, said, slavery, vulnerable, workers, according, trafficking, forced violence, said, harassment, abuse, report, based, sexual, physical
mistral workers, work, forced, labour, children, women, working, conditions sexual, women, harassment, said, abuse, physical, children, workers
mistral post-hoc said, trafficking, forced, labour, slavery, abuse, workers, 2020 sexual, violence, said, abuse, harassment, based, report, women
Figure 5: Rationale token analysis of RaFoLa. (Left) Most frequent tokens extracted from the RaFoLa corpus, human annotations, and rationales extracted from model self-explanations and post-hoc explanations for articles #1 and #8. (Right) Ranking of named entities in articles class #1 extracted from model self-explanations and post-hoc explanations for Llama3. Resulting distributions are compared to the entity distribution found in human annotations.
5 Analyses
In this section, we provide a qualitative analysis of the differences in token selection across languages/risk indicators. We carry out part-of-speech/entity analyses and look into the most frequently selected tokens.
5.1 Sentiment Classification (SST)
POS We apply Part-of-Speech (POS) tagging to the selected tokens for SST/mSST. [^7] We use POS-tagging models from spaCy. We use human annotations as a baseline and analyse the difference in relative distribution over POS tags between the human baseline and self-explanations as well as LRP post-hoc rationales. Results for Mistral are shown in Figure 4, positive values signify relatively higher frequency than in human annotations.
The most important tokens include adjectives (ADJ), nouns (NOUN) and verbs (VERB) followed by adverbs (ADV). Both ADJ and ADV appear much less among selected tokens for both LRP as well as self-explanations in comparison to humans (negative values), with an exception for the English self-explanations (positive values). For Danish and Italian, ADJ/ADV appear even 10%/5% less, in post-hoc LRP than in the human annotations. On the other hand, pronouns (PRON) seem to be more relevant in LRP for English compared to human rationales. [^8] Please note, that we report percentage points here. We further see that nouns are used more across all languages in self-explanations in comparison to human annotations.
For the mSST dataset we see an increased usage of coordinating conjunction (CCONJ), i.e., words like and, but, for in the human annotations which then decreases across all languages for the models. This is most likely an effect of the difference in annotation guidelines for SST and mSST. [^9] Annotators in the SST study were asked to annotate individual words whereas guidelines for mSST required rather precise and connected phrases than individual tokens.
For Llama3 (Figure 12 in the Appendix), we see overall similar results with the main difference in the selection of adjectives in self-explanations. Across all languages we see an increase in the selection of adjectives for self-explanations in comparison to human rationales by about 5%, while the decrease for LRP is similar to the one in Mistral.
Selected Tokens We extract the top-8 most frequent tokens in the SST/mSST corpus. They are shown in Table 4 in the Appendix. Although the significance of those tokens is limited, we can see various trends that support the findings from our POS analysis in the previous section. Across models, we see that rationales in Mistral contain less adjectives and adverbs than for Llama2 and Llama3 in both generated and post-hoc rationales. We also overall see that across all models post-hoc rationales include less adjectives and adverbs than generated self-explanations where Llama2 relies the most on adjectives and adverbs.
5.2 Forced Labour detection (RaFoLa)
Selected tokens
To better understand the differences between model-based rationales and human annotations in detecting indicators of forced labor, we present in Figure 5 (left) the most frequent tokens for indicators #1 and #8, which show the lowest and highest agreements (see Section 4.2). We show other indicators in Table 3 in the Appendix.
When comparing the most frequent tokens as selected by human annotators and different models, we overall find that there is agreement in commonly used tokens, e.g. descriptive nouns covering the general topic of the dataset (workers, work, labour, etc.). Reflecting the shared use of natural language, human annotations and self-explanations show greater overlap in POS selection compared to post-hoc rationales (Figure 13 in the Appendix).
Comparing the indicators, article #8 (Physical and sexual violence) deviates more from the corpus’s most frequent tokens (first row) than #1. Keywords for #8, such as sexual and women, are more clearly identifiable, whereas keywords for #1 (Abuse of Vulnerability) may be harder to detect, possibly due to a less clear definition for an untrained model. This supports our finding that model probability scores are more affected by masking rationales in #8 than for #1 (Section 4.3).
Named Entities
By examining the distribution of named entities across rationale types, we next aim to uncover how annotations extracted from self-explanations and post-hoc explanations differ from humans rationales. Entities are extracted using spaCy and manually verified for correctness. In Figure 5 (right), we show the ranking of named entities across model-based rationales in Llama3 for article #1. Positive/negative scores indicate higher/lower occurrence of a specific entity type as compared to the entity distribution extracted from human annotations. We find that model-based rationales exhibit a stronger emphasis on organizations (ORG) than human rationales ($10-15\%$ more). Compared to models, human rationales instead prioritize broader societal categories, with increased usage of tokens referring to nationalities, religious or political groups (NORP, around $+5\%$), and geopolitical entities (GPE, $+3-8\%$). Post-hoc rationales consistently identify a greater number of person names (PER, $+4\%$) than both humans and self-explanations, as also shown in additional analyses in Appendix E.3 across models and indicators.
6 Discussion
In this paper, we evaluated self-explanations, i.e., explanations generated by instruction-tuned LLMs, based on their plausibility to humans and their faithfulness to models. We instructed 4 LLMs: Llama2, Llama3, Mistral and Mixtral for 2 text classification tasks in English but also in Italian and Danish. We constrained self-explanations to the input tokens of the respective text samples for which established evaluation methods can be easily applied. We analysed the sentiment classification (SST/mSST) and forced labour classification (RaFoLa) for which human annotations are available and included post-hoc feature attribution with layer-wise relevance propagation (LRP) for comparison.
Pairwise comparison between the 3 different types of explanations (humans, generated, post-hoc), i.e., plausibility, shows that human annotations and generated explanations agree much more strongly than post-hoc explanations with either. We further observe that Llama3 shows the highest level of agreement across both tasks and all languages. Our POS analysis further reveals a similar pattern of selecting rationales in SST/mSST between humans and self-explanations where both are mostly selecting adjectives, nouns and verbs. Across tasks, we see more pronounced differences for LRP, selecting less adjectives and adverbs for SST and more nouns in forced labour detection than humans. Fluctuations in human-model agreement for different POS tags, also across languages, have been observed before Brandl and Hollenstein (2022). The distribution of named entities in RaFoLa annotations reveals differences between models and humans, with models emphasizing organizations and person names, while humans annotate more geopolitical and religious entities, suggesting distinct approaches to collecting and evaluating evidence in forced labor detection.
Zero-shot performance in detecting forced labor varies across indicator types, with higher accuracies also resulting in higher agreements to human rationales for indicators #5 and #8. In comparison, the ability of models to solve the sentiment classification task in unseen languages like Danish is remarkable. Not only are the models able to understand the Danish prompt, they also return specific input tokens that are aligned with human rationales. We found that all models were able to make accurate predictions on both Danish and Italian, though it remains unclear whether this is due to data contamination during pre-training, e.g., training data for Llama2 has been reported to contain $0.11\%$ Italian Touvron et al. (2023).
Faithfulness is a key desired property when evaluating the reliability of XAI methods. In our analyses, we found comparable faithfulness scores for self-explanations and post-hoc LRP that suggest that self-explanations could be a more accessible alternative to computation-intensive post-hoc explanations. Besides established post-hoc attribution approaches, the ability of language models to provide self-explanations has offered a direct and human-understandable communication between user and model. This not only enhances usability in particular for lay people but, as presented here, also results in explanations that are similarly faithful and align more closely with human rationales compared to binarized post-hoc attributions. Yet, the generation process behind self-explanations remains obfuscated and may suffer from counterfactuality, enabling the model to give untruthful explanations for correct predictions Ji et al. (2023). In particular, predictions and explanations made by LLMs can identify alternative solutions to task-solving that may not be intuitive to humans. Therefore, good explanations should highlight the learned prediction strategy, ensuring it is faithful to the model’s approach, even if it is not directly plausible Agarwal et al. (2024).
In summary, we find higher plausibility for self-explanations compared to post-hoc rationales while maintaining faithfulness. However, for more challenging tasks, we still see room to improve the plausibility of self-explanations. We further see that LLMs require careful instructions to generate useful self-explanations. We currently do not know why self-explanations align more closely with human annotations than post-hoc attributions. While training procedures such as reinforcement learning from human feedback Ziegler et al. (2019) may incentivize more human-like explanations Agarwal et al. (2024), limited access to models, procedures and data, restricts detailed analysis.
Our study represents a first step toward understanding and building more intuitive model explanations by directly comparing human annotations with those generated by models. Evaluating free-text explanations for factuality, usability, and faithfulness is crucial for ensuring their practical and intuitive application, especially given the growing use of increasingly complex LLMs by lay people who may not fully understand their mechanisms.
Limitations
We acknowledge that annotations may be affected by annotator bias, varying guidelines, and differing expertise, impacting the consistency of rationales. Also the number of annotators and the level of details in the instructions varied across the annotation studies we have considered for this paper. Furthermore, for the forced labour detection, annotations by legal scholars might differ from the ones provided and would also be interesting to compare with model rationales.
We focus our study on rationales based on the input while free text explanations might provide more useful information and pose the more realistic scenario.
While agreement between human and model rationales may be desired, it has been shown in previous work, that humans do not necessarily prefer human-written explanations in comparison to the ones generated by LLMs in the case of free text explanations Wiegreffe et al. (2022).
The high zero-shot performance, especially with SST, may be an effect of data contamination, which is likely part of the training data. We can further not exclude the possibility that rationales or task explanations have been included in the training corpus.
Acknowledgments
We thank our colleagues at the CoAStaL NLP group for constructive feedback in the beginning of the project. In particular, we would like to thank Alice Schiavone and Anders Søgaard for helping us with the prompt translations.
SB received funding by the European Union under the Grant Agreement no. 10106555, FairER. OE is funded by the German Ministry for Education and Research (under refs 01IS18056A and 01IS18025A) and BIFOLD. Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or European Research Executive Agency (REA). Neither the European Union nor REA can be held responsible for them.
Appendix A Instructions
A.1 SST
Figure 6: Prompts in all 3 languages to solve sentiment classification.
Figure 7: Follow-up prompts in all 3 languages to extract rationales.
A.2 RaFoLa
Figure 8: Prompts for classification and rationale extraction for the RaFoLa dataset.
Figure 9: Indicators defined by the International Labour Organization and published by Mendez Guzman et al..
Appendix B F1-Plausibility scores
SST
mSST      English
Danish
Italian
Figure 10: Pair-wise F1 comparison scores between rationales on SST and multilingual SST (English, Danish and Italian). We compare rationales annotated by humans, generated by models, and computed post-hoc with LRP.
RaFoLa       #1 Abuse of vulnerability
#2 Abusive working and living conditions
#5 Excessive Overtime
#8 Physical and sexual violence
Figure 11: Pair-wise F1 comparison scores between rationales on RaFoLa.
Table 2: Ratios of identified rationale tokens by humans and those of self-generated rationales across models for SST and RaFoLa.
ratio ratio model
human llama llama3 mistral mixtral
SST 0.29 0.22 0.25 0.32 0.14
mSST (EN) 0.33 0.23 0.27 0.32 0.15
mSST (DK) 0.38 0.18 0.29 0.40 0.21
mSST (IT) 0.37 0.26 0.31 0.39 0.25
RaFoLA #1 0.05 0.05 0.23 0.12 0.16
RaFoLA #2 0.06 0.04 0.21 0.11 0.12
RaFoLA #5 0.04 0.02 0.08 0.06 0.05
RaFoLA #8 0.07 0.04 0.12 0.08 0.10
Appendix C Rationale ratios
Table 2 presents ratios of identified rationale tokens for SST, multilingual SST (mSST), and RaFoLa datasets in humans and models. The ratio of identified rationale tokens by humans tends to be higher for the shorter movie reviews in SST and mSST, and lower for the longer paragraphs in RaFoLa, suggesting that the human annotators have more sparsely identified relevant phrases when presented with longer texts. This can be attributed to the differences in samples across the datasets. While movie reviews in SST tend to be short and were written with the purpose of expressing a sentiment, the articles in RaFoLa were more descriptive of a specific situation or incident and might or might not include the violation of one (or more) of the 11 risk indicators.
Figure 12: Distribution of POS-tags in comparison to top-6 POS in human annotations for SST. Absolute POS tags have been normalized (so they sum to 1) prior to computing the difference to human baseline. Ranking is based on prevalence in human rationales. Difference is shown for self-explanations (left) and post-hoc LRP (right) across all data splits and languages for Llama3.
Figure 13: Distribution of POS-tags in comparison to top-6 POS in human annotations for RaFoLa. Absolute POS tags have been normalized (so they sum to 1) prior to computing the difference to human baseline. Ranking is based on prevalence in human rationales. Difference is shown for self-explanations and post-hoc LRP across all data splits and languages for Mistral, (left) and Llama3 (right).
Appendix D POS analysis
Figure 12 shows difference in relative POS distribution between model rationales and humans for Llama3 on SST/mSST. Positive values signify higher frequency in comparison to human rationales. Results for this and a similar figure for Mistral (Figure 4) are discussed in Section 5.1. Figure 13 show the same POS analysis for RaFoLa.
Appendix E Most frequent rationale tokens
Table 3: List of top-8 most frequent tokens in the RaFoLa corpus (first row) together with the most frequent rationales as identified by human annotators, as well as self-generated and post-hoc explanations.
#1 Abuse of vulnerability #2 Abusive working and living cond. #5 Excessive overtime #8 Physical and sexual violence
corpus said, workers, labour, human, work, forced, rights, slavery
human workers, work, forced, women, children, labour, said, exploitation workers, conditions, work, forced, water, little, said, working hours, work, week, days, long, claimed, worked, like sexual, abuse, harassment, women, violence, said, verbal, physical
llama2 workers, work, migrant, women, exploitation, forced, labour, children workers, work, conditions, forced, day, working, children, water pay, work, ahmad, received, little, breaks, \$, 600 abuse, sexual, women, harassment, retaliation, verbal, physical, advances
llama2 post-hoc trafficking, world, china, children, said, forced, child, exploitation trafficking, forced, said, world, children, conditions, labour, covid-19 covid-19, workers, thailand, thai, kingdom, california, hours, basi said, violence, women, based, walmart, global, guardian, jennifer
llama3 workers, work, said, labour, forced, women, children, working workers, work, said, conditions, working, labour, forced, day hours, working, day, work, pay, said, workers, days said, women, sexual, abuse, harassment, workers, violence, physical
llama3 post-hoc labour, said, slavery, vulnerable, workers, according, trafficking, forced said, workers, labour, according, conditions, slavery, work, trafficking hours, working, work, overtime, day, workers, forced, said violence, said, harassment, abuse, report, based, sexual, physical
mistral workers, work, forced, labour, children, women, working, conditions workers, work, said, forced, conditions, working, labour, children said, day, mr, work, hours, days, delivery, home sexual, women, harassment, said, abuse, physical, children, workers
mistral post-hoc said, trafficking, forced, labour, slavery, abuse, workers, 2020 said, covid-19, 2019, workers, trafficking, labour, forced, world employer, covid-19, said, years, police, cotton, mr, paying sexual, violence, said, abuse, harassment, based, report, women
mixtral workers, work, said, forced, children, labour, women, abuse workers, work, said, conditions, forced, labour, working, paid hours, work, day, forced, workers, working, days, week sexual, women, said, workers, harassment, abuse, violence, reported
Table 4: List of top-8 most frequent tokens in the SST and mSST corpus together with the most frequent rationales as identified by human annotators, as well as self-generated and post-hoc explanations.
SST mSST English mSST Danish mSST Italian
full corpus movie, film, like, comedy, -, characters, work, romantic film, movie, characters, bad, like, performances, funny, story film, filmen, karakterer, bare, sjov, ‘, se, præstationer film, i, divertente, personaggi, interpretazioni, trama, avvincente, commedia
human funny, movie, beautifully, bad, best, hilarious, stupid, wonderful bad, performances, funny, good, dull, film, compelling, dumb film, sjov, overbevisende, plot, vittig, bare, dårligt, præstationer divertente, avvincente, noioso, interpretazioni, ben, brutto, film, intelligente
llama2 bad, beautifully, best, fun, stupid, wonderful, funny, worst bad, funny, dull, compelling, witty, good, long, little sjov, spændende, underholdende, præstationer, bedste, overbevisende, spænding, vittig divertente, film, avvincente, noioso, interpretazioni, ben, intelligente, senso
llama2 post-hoc movie, comedy, film, little, like, far, funny, stupid movie, film, bad, funny, dull, little, dumb, compelling film, filmen, ‘, sjov, præstationer, karakterer, dårlig, dårligt film, divertente, noioso, avvincente, senso, umorismo, brutto, i
llama3 bad, best, beautifully, compelling, funny, love, little, hilarious funny, bad, performances, dull, compelling, good, little, best sjov, bedste, overbevisende, humor, dårlig, dårligt, kedelig, spænding divertente, avvincente, noioso, umorismo, senso, brutto, intelligente, spiritoso
llama3 post-hoc comedy, like, far, bad, beautifully, best, little, love comedy, like, good, little, big, high, bad, dull ¦, pã¥, vã, film, sã¥, re, spã, sjov film, divertente, cosã¬, avvincente, piã¹, noioso, interpretazioni, brutto
mistral bad, funny, comedy, best, beautifully, compelling, love, far bad, performances, funny, characters, dull, compelling, good, little sjov, filmen, præstationer, film, overbevisende, sjovt, spænding, humor film, divertente, personaggi, interpretazioni, avvincente, noioso, trama, brutto
mistral post-hoc movie, film, bad, -, like, love, feels, year film, bad, funny, dull, good, dumb, characters, comedy film, filmen, filmens, sjov, ‘, præstationer, blanding, spændende film, divertente, interpretazioni, personaggi, commedia, noioso, cinema, avvincente
mixtral best, film, laugh, like, year, love, pretty, good bad, funny, compelling, great, good, witty, performances, intelligent sjov, sjovt, film, præstationer, overbevisende, humor, spændende, tilfredsstillende divertente, film, interpretazioni, noioso, umorismo, i, commedia, trama
llama2 llama3 mistral mixtral
#tokens 0.11 0.04 0.34 0.16
json-syntax 0.86 0.02 0.05 0.37
Table 5: Ratio of mismatches when prompting models for a max. number of tokens and correct json syntax. Averaged scores for SST across all subsets & languages.
E.1 Instruction following
For generating and processing the self-explanations, we instructed the models to return rationales in a json format. Since many of those outputs resulted in SyntaxErrors, we included a syntax check based on Llama3 where we instructed the model in a separate step to correct the json syntax in case such an error occurred. The ability to return correct syntax varied across models. We also saw differences when following the instruction of returning the correct number of rationales for which we set an upper bound for SST based on the human annotations. We show results for SST for all 4 models, averaged across subsets and languages in Table 5. The results show that Llama2 has a lot of difficulties with respect to json syntax with syntax errors occurring in $86\%$ of the case. Both Llama3 and Mistral have a low error rate with $2\%$ and $5\%$ respectively. At the same time, Mistral returns more than the maximum number of requested rationale tokens in 1 out of 3 instructions where Llama3 follows the instruction in most cases. Analysing and evaluating the ability to follow instructions has previously been discussed in Qin et al. (2024); Zeng et al. (2024).
E.2 Contrastive post-hoc explanations
SST Refer to caption
RaFoLa Refer to caption
Figure 14: Comparing plausibility scores for non-contrastive and contrastive post-hoc approaches using Kappa agreement scores. Left: SST and multilingual SST for English, Danish and Italian (top to bottom rows). Right: RaFoLa for classes #1, #2, #5, #8 (cf. Figure 2).
SST
mSST      English
Danish
Italian
RaFoLa        #1
#2
#5
#8
Figure 15: Faithfulness evaluation for SST and mSST (top row) and RaFoLa (bottom row). The probability after masking the tokens extracted from human rationale annotations, generated model rationales (self-explanations) and post-hoc model attributions is compared across models. Lower probability indicates more faithful identification of rationales.
In a complementary analysis, we test the alignment of model rationales with post-hoc attributions, examining whether there is a difference in the plausibility of contrastive and non-contrastive post-hoc explanations. Prior work has suggested that contrastive explanations are more aligned with human reasoning and are thus considered more valuable for humans to understand the model’s decisions Lipton (1990); Miller (2019); Jacovi et al. (2021).
Comparing Cohen’s Kappa scores shown in Figure 14 suggests that contrastive post-hoc approaches do not generally result in higher plausibility than non-contrastive ones. Similarly, contrastive explanations overall exhibit a similar level of faithfulness as non-contrastive explanations (cf. Figure 15). Although contrastive explanations can be more faithful and plausible in certain cases, such as SST and mSST (English) for Llama3, there is no consistent difference between them. This aligns with previously reported high correlation between them Eberle et al. (2023) and the minor observed performance differences for generating the correct label based on contrastive or non-contrastive approaches Krishna et al. (2023).
E.3 Entity analysis
Figure 16: Entity analysis of Rafola for #1 (top) and #8 (bottom) for llama3 (left) and mistral (right), showing top-8 entities across rationale types (human, model, post-hoc).